OpenStudy (anonymous):

For the system shown below, what are the coordinates of the solution that lies in quadrant I? Write your answer in the form (a,b) without using spaces. 3y^2-5x^2=7 3y^2-x^2=23

OpenStudy (anonymous):

3y^2 - 5x^2 = 7 3y^2 = 7 + 5x^2 y = √[(7 + 5x^2)/3] 3y^2 - x^2 = 23 3√[(7 + 5x^2)/3]^2 - x^2 = 23 (7 + 5x^2) - x^2 = 23 4x^2 = 16 x^2 = 4 x = -2 [Not +2 because in the second quadrant, x is negative] y = √[(7 + 5x^2)/3] y = √[(7 + 5(-2)^2)/3] y = √[(7 + 20)/3] y = √9 y = +3 [Not -3 because in the 2nd quadrant y is positive] (-2,3)

OpenStudy (anonymous):

ah its (2,3) it said :/ but thanks !